OpenStudy (anonymous):

Question what is 449.1666667 as a fraction?

OpenStudy (anonymous):

Firstly, we divide 1 by the decimal part: 1/0.166666 = 6. So we know this is the same as \[449 \frac16 = \frac{449*6 +1}{6} = \frac{2695}{6} \]